Suppose X and Y are random variables with P(X = 1) = P(X = -1) = 12; P(Y = 1) = P(Y = -1) = 1
Let c = P(X = 1 and Y = 1).
(a) Determine the joint distribution of X and Y, Cov(X, Y), and Cor(X, Y).
(b) For what value(s) of c are X and Y independent? For what value(s) of c are X and Y 100% correlated?
1
Expert's answer
2020-04-30T15:38:20-0400

"a) P\{X=1, Y=1\}=\frac{1}{3}\\
P\{X=1, Y=-1\}=\frac{1}{6}\\
P\{X=-1, Y=1\}=\frac{1}{6}\\
P\{X=-1, Y=-1\}=\frac{1}{3}\\
COV(X,Y)=M(XY)-M(X)M(Y)\\
M(X)=M(Y)=0\\
COV(X,Y)=M(XY)\\
P\{XY=1\}=\frac{2}{3}\\
P\{XY=-1\}=\frac{1}{3}\\
M(XY)=\frac{1}{3}\\
COV(X,Y)=\frac{1}{3}\\
Cor(X,Y)=\frac{COV(X,Y)}{\sqrt{D(X)D(Y)}}\\
D(X)=M(X^2)-(M(X))^2\\
M(X)=0\\
M(X^2)=1\\
D(X)=1\\
D(Y)=1\\
Cor(X,Y)=COV(X,Y)\\
Cor(X,Y)=\frac{1}{3}\\
b)1)c=P\{X=1, Y=1\}=P\{X=1\}P\{Y=1\}=\frac{1}{4}\\
X \text{ and } Y \text{ are independent for }c=1/4.\\
2)\text{Let }Cor(X,Y)=1\\
Cor(X,Y)=\frac{COV(X,Y)}{\sqrt{D(X)D(Y)}}\\
Cor(X,Y)=COV(X,Y)\\
COV(X,Y)=M(XY)-M(X)M(Y)\\
M(XY)=1\\
\text{Hence } P\{X=1,Y=1\}+P\{X=-1,Y=-1\}=1,\\
P\{X=1,Y=-1\}+P\{X=-1,Y=1\}=0.\\
P\{X=1,Y=1\}=\frac{1}{2}\\
P\{X=-1,Y=-1\}=\frac{1}{2}\\
P\{X=1,Y=-1\}=0\\
P\{X=-1,Y=1\}=0\\
\text{So for } c=\frac{1}{2} \text{ X and Y are 100 percent correlated}."